OpenStudy (anonymous):

Suppose f and g are continuous functions such that g(5) = 3 and lim x → 5 [3f(x) + f(x)g(x)] = 18. Find f(5).

jimthompson5910 (jim_thompson5910):

Hint: You can break up the limit to go from \[\Large \lim_{x\to 5}(3f(x)+f(x)g(x)) = 18\] to \[\Large 3\lim_{x\to 5}(f(x))+\lim_{x\to 5}(f(x))*\lim_{x\to 5}(g(x)) = 18\]

OpenStudy (ikram002p):

lim x → 5 [3f(x) + f(x)g(x)] = 18 hint when a function is continues then \( \lim _{x→x_0}=f(x_0)\) also u can use properties of limits 1_ \( \lim f\pm g =\lim f \pm \lim g\) 2_ \( \lim k f =k\lim f \) s.t k is real number